Using common LISP, write a program that will calculate the area of a regular hexagon (polygon that is equilateral).

will rate ASAP! thank you!
Answer #1

if the answer helped you please UPVOTE and if you have any doubts please COMMENT I will surely help.

The area of a regular hexagon with side length a is  \frac{3\sqrt{3}}{2} a^2 = 2.599 \times a^2

Code:

php6t34MQ.png

Output:

phpZ9DqOE.png

if the answer helped you please upvote and if you have any doubts please comment i will surely help.

Code:

; define a function HexagonArea to find the area of a hexagon
(defun HexagonArea()

; Take input from the user for the side length
; and set that to a variable called side
(princ "Enter side length of the hexagon: ")
(setq side (read))

; calculate area as 2.599 * area * area
(setq area (* 2.599 side side))

; Now print the area of the hexagon
(princ "Area of regular hexagon: ")
(write area))

; call the function to calculate the area
(HexagonArea)